OpenStudy (anonymous):

Find sin θ if cot θ = -2 and cos θ < 0

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

cot θ = -2 and cos θ < 0 cot θ = cosθ/sinθ cosθ=-2sinθ \[\cos θ=\sqrt{1-\sin^2θ}\] \[1-\sin^2 \theta= 4\sin^2 \theta\] \[1/5=\sin^2 \theta\] \[1/\sqrt5= \sin \theta\] \[\sin \theta =\pm 1/\sqrt 5\]

OpenStudy (anonymous):

thats not an option -.-

OpenStudy (anonymous):

its -1/2 -5 \/5/2 or \/5/5

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

@logynklode Sorry this is only way i could think of are the options correct

OpenStudy (anonymous):

yeah. theyre correct.